OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions. Through (−1, −3); perpendicular to the line 2x + 7y + 8 = 0

OpenStudy (anonymous):

It's incorrect

OpenStudy (mimi_x3):

Hm whats the correct answer then so i can work it out

OpenStudy (anonymous):

IDK... i'm trying to figure it out also.

OpenStudy (anonymous):

I got 7x-3y+1=0, but that didn't work.

OpenStudy (mimi_x3):

my bad i forgot that its perpendicular m = 7/2 y = mx + b -3 = 7/2(-1) + b -3 = -7/2 + b b = 2 y = 7/2x + 2 Is it right now ?

OpenStudy (mimi_x3):

y = 7/2x + 1/2 my bad

OpenStudy (anonymous):

In a practice problem they show, the answer ='s to 0

OpenStudy (mimi_x3):

oh in general form ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I get to y-(-3)=7/2 (x-(-1)) Then I'm lost on what to do.

OpenStudy (anonymous):

using the y-y1=m(x-x1) formula

OpenStudy (mimi_x3):

well, i got 2x - 7y - 23 = 0 Not sure if im right though since im not good at these types of questions

OpenStudy (anonymous):

No :(

OpenStudy (mimi_x3):

7x - 2y - 7 = 0

OpenStudy (anonymous):

y-(-3) = 7/2 (x-(-1)) -> 2y-7x-1=0